# Chance Constrained PDE-Constrained Optimal Design Strategies Under High-Dimensional Uncertainty

### v1

@@ Abstract
This study presents an advanced computational framework for the optimal design of thermal insulation components in buildings, utilizing silica aerogel porous materials. The framework aims to achieve superior thermal insulation while maintaining structural integrity of the component under stress concentrations. A multiphase continuum model is employed to simulate the thermomechanical behavior of the insulation system, governed by a set of coupled partial differential equations (PDEs). The design process explicitly accounts for spatially correlated uncertainties in the design parameters, particularly the spatially varying aerogel porosity, resulting in a high-dimensional, PDE-constrained optimization under uncertainty. The optimization problem is formulated using a risk-averse cost functional to balance insulation performance with uncertainty mitigation, incorporating statistical moments of the design objective. Additionally, chance constraints are imposed to limit the probability of stress exceeding critical thresholds. To address the challenges arising from high-dimensional parameters, the optimization leverages a second-order Taylor expansion of both the design objective and the chance constraint functions, combined with a low-rank approximation of the Hessian matrix for efficient evaluation of the generalized eigenvalue problem. This approach supports scalable computations, either directly or as a variance-reduction control variate for Monte Carlo estimations. Combined with a gradient-based optimization approach, we achieve a scalable solution algorithm with dimension-independent computational costs in terms of number of PDE solved. Two- and three-dimensional numerical experiments on the design of thermal breaks in buildings showcase the features of the proposed design under uncertainty framework with respect to accuracy, scalability, and computational cost.
@@
Thermal insulation , Optimal design , PDE-constrained optimization , chance constraint , high-dimensional parameters , uncertainty quantification
journal: International Journal for Numerical Methods in Engineering –
@@ 1 Introduction
Silica aerogels have emerged as a premier class of porous materials for high-performance thermal insulation, widely utilized across diverse engineering applications due to their exceptionally low thermal conductivity [1, 2, 3]. Among these applications, aerogel-based insulation components for building envelopes hold significant promise, offering the potential to enhance energy efficiency, reduce carbon emissions, and promote environmental sustainability. However, the adoption of silica aerogels in such systems is hindered by their inherently low mechanical strength and susceptibility to failure under localized stress concentrations. Recent advancements in extrusion-based additive manufacturing may be able to address these limitations by enabling precise spatial control over material deposition, thereby enhancing the mechanical stability of components. These developments have paved the way for integrating silica aerogels into lightweight thermal break insulation elements [2, 4], which effectively interrupt heat transfer pathways and mitigate the adverse effects of thermal bridging at assembly interfaces [5, 6]. Nevertheless, significant uncertainties remain regarding the thermal and mechanical performance of additively manufactured aerogel components, posing challenges to their adoption as reliable and sustainable solutions for thermal break applications in building envelopes.
In recent years, significant efforts have been dedicated to developing computational methods for topology optimization governed by partial differential equations (PDEs) (e.g., [7, 8, 9, 10, 11, 12, 13, 14, 15]). However, the optimal design of additively manufactured aerogel thermal breaks poses two major computational challenges. The first challenge arises from uncertainties in design parameters, stemming from variations in material properties, external loads, and environmental conditions. In particular, the layer-by-layer manufacturing process of aerogel components introduces spatially correlated uncertainties along the layer deposition direction, necessitating the representation of these uncertainties as spatially varying random fields. When discretized, the resulting high-dimensional parameter space makes the optimization problem computationally prohibitive due to the “curse of dimensionality” and the need for repeated evaluations of nonlinear PDEs governing material behavior. The second challenge involves incorporating stress constraints into the optimization process, which introduces nonlinearity and numerical instability. Stress gradients often exhibit discontinuities, further complicating the numerical implementation. To mitigate these issues, global stress constraints that aggregate stress measures (e.g., von Mises stress) across the design domain are commonly employed, significantly reducing the number of constraints. However, selecting an efficient stress aggregation function with appropriate parameters is crucial to achieving computational efficiency while ensuring the detection of localized stress concentrations. Furthermore, stress measures are influenced by uncertainties in material properties and operating conditions, necessitating the incorporation of randomness into stress constraints during the optimization process.
@@
To address uncertainty in optimization, one approach is to utilize a risk-averse formulation of the cost functional, that takes into account the statistical moments of the objective, e.g., [16, 17, 18, 19, 20, 21, 22, 11, 23, 24]. However, these methods face significant computational challenges when evaluating the statistical properties of objectives or conditional probabilities, especially in the context of PDE-constrained optimization within high-dimensional parameter spaces. Various approximations have been proposed to address these challenges. Ghattas et al. [25, 20, 19, 18] utilized Taylor expansions for scalable, dimension-independent solutions with significant computational savings. This linearization, previously applied in robust design [26, 27] and topology optimization [28], was extended by Kriegesmann et al. [21, 29, 30] with a first-order second-moment method for efficient variance gradient evaluation. Tan et al. [31] extended the Taylor approximation method to a multi-objective PDE-constrained optimization framework for multi-material systems under uncertainty. Their approach incorporates phase-field regularization to effectively manage spatial sparsity and the material interface. Chen and Royset [32] provided further analysis of approximation-based algorithms for PDE-constrained optimization under uncertainty. In the context of stress constraints, several studies have investigated the integration of these constraints into optimization frameworks for diverse materials and manufacturing processes [33, 34, 35, 36, 37, 14, 38, 39, 40, 41]. Significant efforts have also focused on incorporating uncertainty into stress-constrained optimization [42], addressing factors such as variable loading and heat sources. For instance, Silva et al. [35] developed a method to account for uncertainties in boundary conditions during manufacturing. Similarly, Oh et al. [43] proposed a topology optimization approach that incorporates uncertainties in loading positions.
Building on previous work, this study presents a scalable computational framework for risk-averse PDE-constrained optimization under spatially correlated and high-dimensional uncertainty, while incorporating chance constraints. The framework is applied to the design of thermal breaks in building envelope, aiming to maximize thermal insulation performance and ensure mechanical stability by mitigating stress concentrations through chance constraints. The forward model is based on a thermomechanical continuum mixture theory for porous materials. The PDE model parameters, such as solid and fluid conductivities and solid elastic modulus, have been determined through model calibration in prior studies. Material porosity, modeled as a spatially correlated random field, serves as the design parameter. Discretization of this field transforms the problem into a high-dimensional optimization under uncertainty To achieve target performance levels while mitigating the uncertainty, the framework employs a risk-averse formulation that integrates the mean and variance of the design objective. The computational burden of Monte Carlo moment estimation is mitigated using second-order Taylor approximations for the design objective and chance constraints. An inexact Newton-Conjugate Gradient algorithm with convergence properties independent of problem dimension is utilized to ensure computational scalability—measured by the number of PDE solutions—regardless of the dimensionality of the design parameter space. The effectiveness and computational cost of the framework are explored through two- and three-dimensional numerical examples, including a case study of a beam-thermal break component in the building envelope. Compared to our earlier work in [31], the key contributions of this study are as follows: (1) Incorporation of probabilistic constraints in the optimization framework to mitigate stress concentrations. (2) Utilization of second-order Taylor approximations as a control variate for Monte Carlo estimation of the design objective, enabling more accurate optimal designs in high-variance uncertainty scenarios. (3) Implementation of an inexact Newton-Conjugate Gradient algorithm, facilitating parallelization and accelerating PDE-constrained optimization calculations. (4) Extension of the numerical framework to three-dimensional design of insulation components.
The rest of the paper is organized as follows: Section 2 presents the optimal design under uncertainty problem, including the thermomechanical forward model, the modeling of spatially correlated uncertainty in the design parameter, and the formulation of mean-variance optimization. Section 3 is dedicated to the approximation methods, encompassing Taylor approximation and the Monte Carlo estimate with control variate of the moments of the design objective. Section 4 describes the gradient-based optimization method and explains the computation of the gradient of the cost functional concerning the design parameter. Section 5 presents the numerical experiments on the optimal design of thermal breaks, followed by the conclusions and future works in Section 6.
@@ 2 Problem formulation: Design of aerogel thermal break under uncertainty
This study investigates the optimal design of building insulation components using silica aerogel materials. Figure 1 illustrates the schematic representation of a building thermal break, consisting of a concrete beam integrated with an aerogel insulation system, along with the thermal and mechanical boundary conditions employed in the numerical experiments. Concrete beams significantly influence the thermal performance of building envelopes due to their high thermal conductivity, which creates preferential pathways for heat transfer. The inclusion of silica aerogel thermal breaks disrupts this conduction, effectively mitigating thermal bridging from the interior to the exterior of the building envelope. However, thermal expansion of the concrete beam can induce stress concentrations within the silica aerogel insulation component, potentially leading to mechanical failure and compromising the thermal break’s effectiveness. In this section, we introduce the problem of optimal design under uncertainty, leveraging a forward model of aerogel and concrete materials governed by a system of partial differential equations (PDEs). We then describe how the uncertainty in design parameters are modeled, as well as formulation of design problem.
@@
Figure 1: The domain of a beam-insulator system used for numerical experiments, indicating both the mechanical (left) and thermal (right) boundary conditions.
2.1 Forward thermomechanical model and design objectives
For silica aerogel materials, we utilize a thermomechanical model based on continuum mixture theory [44]. The model accounts for both an incompressible solid aerogel phase and a compressible fluid (gaseous) phase within the material. According to the mixture theory, each phase independently adheres to its own motion, balance laws, internal energy, and entropy, while collectively following the overall balance laws governing the entire mixture. As a result, a system of four partial differential equations is derived, which governs the heat transfer in the solid and fluid phases,and the deformation in the solid phase and the evolution of pore pressure. For our problem of designing thermal breaks, we focus on steady-state and static conditions, which simplify the governing equations to the following PDEs,
@@ 2.1 Forward thermomechanical model and design objectives
$\begin{cases}-\nabla\cdot(\phi_{s}\,\kappa_{s}\,\nabla\theta_{s})+h(\theta_{s}-\theta_{f})=0\quad&\text{in}\quad\Omega,\\ -\nabla\cdot(\phi_{f}\,\kappa_{f}\,\nabla\theta_{f})-h(\theta_{s}-\theta_{f})=0\quad&\text{in}\quad\Omega,\\ -\phi_{s}\,\kappa_{s}\,\nabla\theta_{s}=h_{\mathrm{air}}(\theta_{s}-\theta_{\mathrm{amb}})\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{4},\\ -\phi_{f}\,\kappa_{f}\,\nabla\theta_{f}=h_{\mathrm{air}}(\theta_{f}-\theta_{\mathrm{amb}})\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{4},\\ \end{cases}$ (1)
@@
$\begin{cases}D\,p+(\nabla\cdot\mathbf{u}_{s})=0\quad&\text{in}\quad\Omega,\\ -\nabla\cdot\mathbf{T}^{\prime}_{s}-(2\,\phi_{f}-1)\,\nabla p=0\quad&\text{in}\quad\Omega,\\ \mathbf{u}_{s}=\bar{\mathbf{u}}_{s}\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{3},\\ \nabla\mathbf{u}_{s}.\mathbf{n}=0\quad&\text{on}\quad\Gamma_{2}\end{cases}$ (2)
where model parameters are $\theta=(\kappa_{s},\kappa_{f},D,\mathbf{E}_{s})$ and states are $u=(\theta_{s},\theta_{f},\mathbf{u_{s}},p)$. The state variables $\theta_{s}$, $\theta_{f}$, $\mathbf{u}_{s}$, and $p$ represent the solid and fluid temperatures, solid displacement, and fluid pressure, respectively, defined over the domain $\Omega$. Here, $\phi_{f}(\mathbf{x})$ represents the volume fraction of fluid (porosity), and $\phi_{s}(\mathbf{x})$ represents the volume fraction of solid, which can alternatively be written as $1-\phi_{f}(\mathbf{x})$. The solid displacement is subject to fixed values $\bar{\mathbf{u}}_{s}$ at the boundaries $\Gamma_{1}$ and $\Gamma_{3}$. A roller boundary condition is implemented at $\Gamma_{2}$. At $\Gamma_{1}$ and $\Gamma_{4}$, we consider convective heat transfer with a coefficient $h_{\mathrm{air}}$ and ambient temperature $\theta_{\mathrm{amb}}$. A Neumann type boundary condition with no heat flux is implemented at $\Gamma_{2}$. In the above equations, $\Omega\subset\mathbb{R}^{d}$ is an open and bounded domain with Lipschitz boundary $\partial\Omega$, and $\Gamma_{1},\Gamma_{3},\Gamma\subset\partial\Omega$ denote the Dirichlet and $\Gamma_{2}\subset\partial\Omega$ indicate the Neumann portions of $\partial\Omega$ such that $\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{4}=\partial\Omega$, $\Gamma_{1}\cap\Gamma_{2}=\emptyset$, $\Gamma_{3}\cap\Gamma_{4}=\Gamma_{3}$, and $\Gamma_{1}\cap\Gamma_{4}=\emptyset$. In (2), the effective solid stress is
$\mathbf{T}^{\prime}_{s}=2\,\mu\,\mathbf{E}_{s}+\lambda\,\mathrm{tr}(\mathbf{E}_{s})\,\mathbf{I}\quad\text{with}\quad\mathbf{E}_{s}=\frac{1}{2}\big{(}\nabla\mathbf{u}_{s}+(\nabla\mathbf{u}_{s})^{T}\big{)},$ (3)
where $\mathbf{E}_{s}$ is the solid strain and $\lambda$ and $\mu$ are the Lamé constants. Determining the model parameters $\kappa_{s}$, $\kappa_{f}$, $h$, $K$, $C$, and $\mu$ from experimental data of silica aerogel is demonstrated in [45] and assumed to be known. for the concrete beam we consider the following thermoelastic model,
@@
$\begin{cases}-\nabla\cdot(\,\kappa_{b}\,\nabla\theta_{b})=0\quad&\text{in}\quad\Omega_{b},\\ \kappa_{b}\,\nabla\theta_{b}=h_{\mathrm{air}}(\theta_{b}-\theta_{\mathrm{amb}})\quad&\text{on}\quad\Gamma_{3},\\ \nabla.\mathbf{T}_{b}=0&\text{in}\quad\Omega_{b}\\ \end{cases}$ (4)
where $\kappa_{b}$ is thermal conductivity of beam, $\theta_{b}$ represents temperature of beam and $\mathbf{T}_{b}$ is the beam stress. In (4), the stress in beam is
$\mathbf{T}_{b}=2\mu_{b}\mathbf{E}_{b}+\lambda_{b}\mathrm{tr}(\mathbf{E}_{b})\mathbf{I}-\alpha_{T}(3\lambda_{b}+2\mu_{b})(\theta_{b}-\theta_{0})\quad\text{with}\quad\mathbf{E}_{b}=\frac{1}{2}\big{(}\nabla\mathbf{u}_{b}+(\nabla\mathbf{u}_{b})^{T}\big{)}$ (5)
@@
where $\mathbf{E}_{b}$ is strain in the beam, and $\lambda_{b}$ and $\mu_{b}$ are the Lamé constants for the beam, $\alpha_{T}$ is coefficient of thermal expansion and $\theta_{0}$ is the reference temperature.
In the following sections, the forward state problems (1), (2) and (4) are expressed in an abstract form as,
$\mathcal{R}(\mathbf{u},m,d)=0\quad\mathrm{in\ }\mathcal{V}^{\prime},$ (6)
where, given the uncertain parameter $m\in\mathcal{M}$ and design parameter $d\in\mathcal{D}$, the goal is to find the state $\mathbf{u}=(\theta_{s},\theta_{f},\mathbf{u}_{s},p,\mathbf{u}_{b})\in\mathcal{U}$ such that the PDE residual operator $\mathcal{R}(\cdot)$ is minimized to zero in $\mathcal{V}^{\prime}$. Here $\mathcal{M}$ and $\mathcal{D}$ represent infinite-dimensional Hilbert spaces, and $\mathcal{V}^{\prime}$ denotes the corresponding dual space.
Weak formulation
To derive the weak formulation of the governing equations, we start with the partial differential equations (PDEs) (1) ,(2) and (4). The problem can be stated as follows: Find $\mathbf{u}=(\theta_{s},\theta_{f},\theta_{b},\mathbf{u}_{s},\mathbf{u}_{b},p)\in\mathcal{U}$, such that
$-\big{<}\phi_{s}\kappa_{s}\nabla\theta_{s},\nabla z_{s}\big{>}-\big{<}h(\theta_{s}-\theta_{f}),z_{s}\big{>}-\big{<}\phi_{s}h_{\mathrm{air}}(\theta_{s}-\theta_{\mathrm{amb}}),z_{s}\big{>}$ $=$ $0,$
$-\big{<}\phi_{f}\kappa_{f}\nabla\theta_{f},\nabla z_{f}\big{>}+\big{<}h(\theta_{f}-\theta_{f}),z_{f}\big{>}-\big{<}\phi_{f}h_{\mathrm{air}}(\theta_{f}-\theta_{\mathrm{amb}}),z_{f}\big{>}$ $=$ $0,$
$-\big{<}\kappa_{b}\nabla\theta_{b},\nabla z_{b}\big{>}-\big{<}h_{\mathrm{air}}(\theta_{b}-\theta_{\mathrm{amb}}),z_{b}\big{>}$ $=$ $0,$
@@
$-\big{<}(\lambda_{b})(\nabla\cdot\mathbf{u}_{b})\mathbf{I}+\mu_{b}\big{(}\nabla\mathbf{u}_{b}+(\nabla\mathbf{u}_{b})^{T}\big{)},\boldsymbol{w}_{b}\big{>}+\big{<}\mathbf{t},\boldsymbol{w}_{b}\big{>}$ $=$ $0,,$

### v2

# Chance-Constrained Optimal Design of Porous Thermal Insulation Systems Under Spatially Correlated Uncertainty

@@ Abstract
This paper presents a computationally efficient method for the optimal design of silica aerogel porous material systems, balancing thermal insulation performance with mechanical stability under stress concentrations. The proposed approach explicitly accounts for additive manufacturing uncertainties by modeling material porosity as a spatially correlated stochastic field within a multiphase finite element formulation. A risk-averse objective function, incorporating statistical moments of the design objective, is employed in conjunction with chance constraints that enforce mechanical stability by restricting the probability of exceeding critical stress thresholds. To mitigate the prohibitively high computational cost associated with the large-dimensional uncertainty space and Monte Carlo estimations of the objective function’s statistical moments, a second-order Taylor expansion is utilized as a control variate. Furthermore, a continuation-based smoothing strategy is introduced to address the non-differentiability of the chance constraints, ensuring compatibility with gradient-based optimization. The resulting framework achieves computational scalability, remaining agnostic to the dimensionality of the stochastic design space. The effectiveness of the method is demonstrated through numerical experiments on two- and three-dimensional thermal break systems for building insulation. The results highlight the framework’s capability to solve large-scale, chance-constrained optimal design problems governed by finite element models with uncertain design parameter spaces reaching dimensions in the hundreds of thousands.
@@
Thermal insulation , Optimal design , chance constraint optimization , high-dimensional parameters , uncertainty quantification
journal: Journal of Risk and Uncertainty in Engineering Systems –
@@ 1 Introduction
The development of high-performance insulation materials is critical for improving energy efficiency and reducing the environmental impact of buildings. Silica aerogels, renowned for their ultra-low thermal conductivity [1, 2, 3], offer exceptional insulating properties but are highly susceptible to mechanical failure due to stress-concentration. Advances in additive manufacturing have facilitated precise control over the spatial distribution of material properties, potentially enabling mechanically robust and lightweight insulation systems. However, the inherent uncertainties in the thermal and mechanical performance of these engineered materials pose significant challenges to their reliable integration into construction applications, necessitating robust design strategies.
While finite element-based topology optimization techniques have advanced considerably [4, 5, 6, 7, 8, 9, 10, 11, 12], design of additively manufactured aerogel components presents two primary challenges. The first challenge arises from uncertainties in spatially varying design parameters, particularly those induced by the layer-by-layer fabrication process. This manufacturing variability introduces spatially correlated stochastic fluctuations along the deposition direction, leading to a high-dimensional design space when discretized using finite element methods. Consequently, the computational cost becomes prohibitive due to the necessity of solving multiple nonlinear finite element solution of the porous material response. The second challenge involves incorporating stress constraints into the optimization framework. These constraints introduce significant nonlinearity and discontinuities, particularly when addressing localized stress concentrations. Global stress constraints, often formulated using aggregated measures such as von Mises stress, require carefully chosen aggregation functions to maintain both computational efficiency and sensitivity to critical stress regions. Furthermore, the presence of uncertainty necessitates a stochastic formulation of stress constraints, integrating probabilistic constrained into the optimal design formulation.
@@
To address the first challenge and ensure robustness in optimal design under uncertainty, risk-averse optimization methods incorporates statistical moments, such as mean and variance, into the objective function [13, 14, 15, 16, 17, 8, 18, 19, 20]. However, directly integrating these moments into finite element-based optimization with high-dimensional parameters imposes significant computational costs. Various approximation techniques have been developed to alleviate this burden. Notably, Ghattas et al. [21, 14, 13] demonstrated that Taylor series expansions of statistical moments provide scalable and computationally efficient solutions for optimization problems governed by finite element models. Kriegesmann et al. [16, 22, 23] introduced first-order second-moment methods for efficient gradient-based variance estimation. More recently, Tan et al. [24] extended Taylor-based approximations to multi-objective optimization, incorporating phase-field regularization to enhance spatial sparsity control. To address the second challenge, several studies have explored the integration of stress constraints into optimization frameworks across various materials and manufacturing processes [25, 26, 27, 28, 29, 11, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Efforts have also focused on incorporating uncertainty into stress-constrained optimization [40], addressing factors such as variable loading and heat sources. For example, Silva et al. [27] accounted for uncertainties in boundary conditions during manufacturing, while Oh et al. [41] incorporated uncertainties in loading positions in topology optimization. Mart’ınez and Ortigosa [42] developed a risk-averse approach for fail-safe structures, treating collapse as an uncertainty source rather than a deterministic event. Liu et al. [43] introduced interval reliability-based topology optimization, considering multi-source uncertainties in dynamic responses.
We present a novel computational framework for the optimal design of aerogel-based insulation systems under spatially varying stochastic design parameters, incorporating chance constraints to ensure mechanical stability. The thermomechanical property of the porous silica aerogels is characterized by a multiphase finite element model, where material porosity is represented as a spatially correlated stochastic field. The objective is to optimize the spatial distribution of material porosity to maximize thermal insulation while mitigating stress concentrations of the insulation component. To alleviate the computational burden associated with Monte Carlo-based moment estimation in the mean-variance cost functional, we employ second-order Taylor approximations to efficiently approximate both the design objective and chance constraints. Building on our previous work in [24], this study introduces the following key contributions: (1) Explicit integration of probabilistic stress constraints to directly regulate stress concentrations within the optimization framework, ensuring mechanical robustness. (2) Rather than directly employing the Taylor expansion of the moments of the design objective, this approximation is leveraged as a control variate in Monte Carlo estimation, substantially enhancing accuracy in high-variance uncertainty scenarios. (3) Implementation of an inexact Newton-Conjugate Gradient algorithm, enabling scalable optimization for large-scale problems while maintaining convergence guarantees. The effectiveness and computational efficiency of the proposed approach are demonstrated through two- and three-dimensional numerical experiments of a beam-thermal break component within a building envelope.
The remainder of the paper involves presenting the thermomechanical model for silica aerogels, incorporating spatially correlated uncertainties and the mean-variance cost functional for optimal design in Section 2. Section 3 presents the Taylor approximations and Monte Carlo estimators with control variates. Section 4 details the integration of chance constraints within the gradient-based optimization framework. Numerical experiments demonstrating the method’s effectiveness are provided in Section 5, followed by conclusions and directions for future research in Section 6.
@@ 2 Problem formulation: Design of aerogel thermal break under uncertainty
Figure 1 illustrates a building insulation system involving a concrete beam with an integrated silica aerogel thermal break. Concrete’s high thermal conductivity promotes heat transfer through building envelopes, forming thermal bridges, while thermal breaks mitigate this effect. However, thermal expansion of the concrete beam can generate stress concentrations in the aerogel, risking mechanical failure and reducing insulation effectiveness. This section formulates the optimal design problem under uncertainty, incorporating thermomechanical models governed by partial differential equations (PDEs) and detailing the modeling of uncertainty in design parameters.
@@
Figure 1: The domain of a beam-insulator system used for numerical experiments, indicating both the mechanical (left) and thermal (right) boundary conditions. The solid displacement are fixed at the boundaries $\Gamma_{1}$ and $\Gamma_{3}$ and roller boundary conditions are implemented at $\Gamma_{2}$. At $\Gamma_{1}$ and $\Gamma_{4}$, we consider convective heat transfer cold ambient temperature in the building exterior and hot ambient temperature in the building interior. no heat flux boundary is implemented at $\Gamma_{2}$.
2.1 Thermomechanical finite element models and design objectives
For silica aerogel materials, we adopt a multiphase model based on continuum mixture theory [44], capturing the interactions between an incompressible solid aerogel phase and a compressible fluid (gaseous) phase. The governing equations consist of a system of PDEs that describe heat transfer in both phases, solid-phase deformation, and pore pressure evolution, over the domain $\Omega$ and boundaries $\Gamma_{1}$, $\Gamma_{2}$, $\Gamma_{3}$, and $\Gamma_{4}$, as
@@ 2.1 Thermomechanical finite element models and design objectives
$\begin{cases}-\nabla\cdot(\phi_{s}\,\kappa_{s}\,\nabla\theta_{s})+h(\theta_{s}-\theta_{f})=0\quad&\text{in}\quad\Omega,\\ -\nabla\cdot(\phi_{f}\,\kappa_{f}\,\nabla\theta_{f})-h(\theta_{s}-\theta_{f})=0\quad&\text{in}\quad\Omega,\\ -\phi_{s}\,\kappa_{s}\,\nabla\theta_{s}=h_{\mathrm{conv}}(\theta_{s}-\theta_{\mathrm{a}})\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{4},\\ -\phi_{f}\,\kappa_{f}\,\nabla\theta_{f}=h_{\mathrm{conv}}(\theta_{f}-\theta_{\mathrm{a}})\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{4},\\ \end{cases}$ (1)
@@
$\begin{cases}D\,p+(\nabla\cdot\mathbf{u}_{s})=0\quad&\text{in}\quad\Omega,\\ -\nabla\cdot\mathbf{T}^{\prime}_{s}-(2\,\phi_{f}-1)\,\nabla p=0\quad&\text{in}\quad\Omega,\\ \mathbf{u}_{s}=\mathbf{0}\quad&\text{on}\quad\Gamma_{1}\cup\Gamma_{3},\\ \nabla\mathbf{u}_{s}\cdot\mathbf{n}=0\quad&\text{on}\quad\Gamma_{2}\end{cases}$ (2)
where $\nabla$ is gradient operator, $\mathbf{n}$ is the outward unit normal vector, and the effective solid stress is $\mathbf{T}^{\prime}_{s}=2\,\mu\,\mathbf{E}_{s}+\lambda\,\mathrm{tr}(\mathbf{E}_{s})\,\mathbf{I}$ with $\mathbf{E}_{s}(\mathbf{u}_{s})$ being the solid strain. The model parameters are solid conductivity $\kappa_{s}$, fluid conductivity $\kappa_{f}$, fluid compressibility $D$, the Lamé constants $\lambda$ and $\mu$, and convective heat coefficient $h_{\mathrm{conv}}$ with respect to hot or cold ambient temperature $\theta_{\mathrm{a}}$, and are assumed to be known for silica aerogel from previous studies. The state variables are $\theta_{s}$, $\theta_{f}$, $\mathbf{u}_{s}$, and $p$ represent the solid and fluid temperatures, solid displacement, and fluid pressure, respectively. Additionally, $\phi_{s}$ and $\phi_{f}$ are the volume fractions of solid and fluid (porosity) phases. The concrete beam is assumed to follow a thermoelastic model,
@@
$\begin{cases}-\nabla\cdot(\,\kappa_{b}\,\nabla\theta_{b})=0\quad&\text{in}\quad\Omega_{b},\\ -\kappa_{b}\,\nabla\theta_{b}=h_{\mathrm{conv}}(\theta_{b}-\theta_{\mathrm{a}})\quad&\text{on}\quad\Gamma_{3},\\ \nabla\cdot\mathbf{T}_{b}=0&\text{in}\quad\Omega_{b}\\ \end{cases}$ (3)
@@
where $\theta_{b}$ represents temperature of beam and Cauchy stress is $\mathbf{T}_{b}=2\mu_{b}\mathbf{E}_{b}+\lambda_{b}\mathrm{tr}(\mathbf{E}_{b})\mathbf{I}-\alpha_{T}(3\lambda_{b}+2\mu_{b})(\theta_{b}-\theta_{0})$ with $\mathbf{E}_{b}$ is strain in the beam. In (3), $\kappa_{b}$ is thermal conductivity of beam, $\lambda_{b}$ and $\mu_{b}$ are the Lamé constants for the beam, $\alpha_{T}$ is coefficient of thermal expansion and $\theta_{0}$ is the reference temperature. Finite element solution of the above model requires the weak formulation of the governing equations as:
Find $\mathbf{u}=(\theta_{s},\theta_{f},\theta_{b},\mathbf{u}_{s},\mathbf{u}_{b},p)$, such that
$-\big{<}\phi_{s}\kappa_{s}\nabla\theta_{s},\nabla z_{s}\big{>}-\big{<}h(\theta_{s}-\theta_{f}),z_{s}\big{>}-\big{<}\phi_{s}h_{\mathrm{conv}}(\theta_{s}-\theta_{\mathrm{a}}),z_{s}\big{>}$ $=$ $0,$
$-\big{<}\phi_{f}\kappa_{f}\nabla\theta_{f},\nabla z_{f}\big{>}+\big{<}h(\theta_{f}-\theta_{f}),z_{f}\big{>}-\big{<}\phi_{f}h_{\mathrm{conv}}(\theta_{f}-\theta_{\mathrm{a}}),z_{f}\big{>}$ $=$ $0,$
$-\big{<}\kappa_{b}\nabla\theta_{b},\nabla z_{b}\big{>}-\big{<}h_{\mathrm{conv}}(\theta_{b}-\theta_{\mathrm{a}}),z_{b}\big{>}$ $=$ $0,$
@@
$-\big{<}(\lambda_{b})(\nabla\cdot\mathbf{u}_{b})\mathbf{I}+\mu_{b}\big{(}\nabla\mathbf{u}_{b}+(\nabla\mathbf{u}_{b})^{T}\big{)},\boldsymbol{w}_{b}\big{>}+\big{<}\mathbf{t},\boldsymbol{w}_{b}\big{>}$ $=$ $0,$ (4)